# The Mixed Birth-death/death-Birth Moran Process

### v1

@@ 1 Introduction
Evolution occurs in populations of reproducing individuals. The modern synthesis in the early $20^{\text{th}}$ century fused natural selection with Mendelian inheritance into a quantitative framework [huxley1942evolution]. Within this framework, population genetics studies how alleles spread through a population, with particular focus on fixation—the event that a mutant allele takes over a resident background [fisher1923xxi, haldane1927mathematical, Wright_1931, moran1959survival, kimura1962probability, kimura1979fixation, kimura1980average, nei1983estimation, charlesworth1997rapid, whitlock2000fixation, przeworski2003estimating, patwa2008fixation, cvijovic2015fate]. Evolutionary graph theory is a general framework to study evolutionary dynamics in structured populations [nowak2003linear, lieberman2005evolutionary, nowak2006evolutionary, Ohtsuki_Hauert_Lieberman_Nowak_2006]. A common class of models studied in evolutionary graph theory is variations of what are referred to as Moran processes. In this model, a population of size $n$ is represented by a (possibly weighted, directed) graph on $n$ locations, where edges constrain who can replace whom [^1] The original Moran process, described by Patrick Moran in 1958 [enwiki:1277295443, moran1958random], considered a well-mixed population where each individual can replace any other individual, which is equivalent to the graph being complete. Generalizations to the process on graphs were introduced early in the development of evolutionary graph theory [lieberman2005evolutionary, nowak2006evolutionary]. We follow the modern literature (e.g. [diaz_approximating_2014, diaz2016absorption, goldberg2020phase]) and use the term “Moran process” to refer to more generalised processes, where additional features such as graph structure and the order of the birth and death steps may also be considered. [lieberman2005evolutionary]. In this work, we consider unweighted, undirected graphs. Two types of individuals live on the graph, with one individual at each vertex: residents, who have fitness $1$, and mutants, who have fitness $r$.
@@
In this paper, we explore a smooth interpolation between the two processes, which we call the $\lambda$-mixed Moran process. We unify these models with a single parameter $\lambda\in[0,1]$; in each step, the process performs a Bd update with probability $\lambda$ and a dB update with probability $1-\lambda$. While seemingly simple, the model exhibits interesting properties. For example, we give an example where the fixation probability for $\lambda\in(0,1)$ is bounded by its values at $\lambda=0$ and $\lambda=1$, implying also that monotonicity in $\lambda$ does not hold.
@@ 2.1 Models and notation
A folklore fact is that the fixation probability is additive when $r=1$. We include a proof in Appendix˜A for completeness.
@@
Consider any graph $G$, when $r=1$, for all $\lambda$,
@@ 2.2 Technical tools
We use a potential-function (drift) method. Fix a potential function $\phi$ on subsets $S\subseteq V$. Our goal is to bound the expected one-step change of $\phi$. These theorems are relatively standard, so we defer the proofs to Appendix˜A.
@@ 3 The case $\lambda=1/2$
The claim follows from Theorem˜2.5 and 2.3. ∎
@@
Because $\phi(S)=0$ at $S=\emptyset$ and $\phi(S)=n$ at $S=V$, applying Theorem˜2.4 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ gives
@@
Recall that the fixation probability with initial mutants $S_{0}$ is $|S_{0}|/n$ when $r=1$. A simple coupling argument shows that the fixation probability is monotone with respect to $r$, yielding the following corollary.
@@
Applying Theorem˜2.6 together with Theorem˜3.3 and Corollary˜3.4 gives an FPRAS for Mixed Moran fixation on any graph when $\lambda=1/2$ and $r\geqslant 1$.
@@ 4 Almost Regular Graphs and Random Graphs
We first show that for a regular graph $G$, the fixation probability starting with an initial mutant set $S_{0}$ when $r=1$ is $|S_{0}|/n$, regardless of $\lambda$ and the position of the mutants. This could be compared with Theorem˜3.1 where the same property holds for $\lambda=1/2$ and all graphs: both cases demonstrate this property because the process balances the change in different directions, resulting in the noted invariance. It is not surprising that we use $\phi(S)=|S|$ to prove this.
@@
For any regular graph $G$, we have $\mathrm{fp}^{\lambda,r=1}_{G}(S_{0})=|S_{0}|/n$ for all $S\subseteq V$ and $\lambda$.
Proof.
Fix $S_{0}=\{u\}$ for an arbitrary vertex $u$. Take the potential $\phi(S)=|S|$. Since all degrees are equal
Thus $\phi(S)$ is a martingale. The theorem follows from Theorem˜2.5 and 2.3. ∎
[TODO: reviewer asks whether for regular graphs, there is a closed-form solution for fixation probability]
@@
Note that without loss of generality $\alpha\geqslant(n-1)/(n-2)$ or $\alpha=1$. The latter is trivial, so assume the former. For $n\geqslant 3$, since $r\geqslant\alpha^{2}$ and $\frac{n-1}{n-2}\geqslant\frac{n}{n-1}$,
@@
Applying Theorem˜2.4 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$.
For fixation probability, apply Theorem˜2.5 with $k_{0}=1$ and $k_{1}=n$ gives $\mathrm{fp}(S_{0})\geqslant\frac{1}{n}$}.
@@
Applying Theorem˜2.4 with $k_{1}=n^{2}$, and $k_{2}=(r-1)/(2rn^{2})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$. For the fixation probability, applying Theorem˜2.5 with $k_{0}\geqslant 1$ and $k_{1}=n^{2}$ gives $\mathrm{fp}(S_{0})\geqslant n^{-2}$}. ∎
@@
Combining Theorem˜2.6, 4.3, and 4.5 yields the FPRAS for random graphs.
@@ 5 Bidegreed Graphs
and $\phi(S_{i})$ is a martingale. Applying theorem˜2.5 with $k_{0}=f(\deg_{u})$ and $k_{1}=\sum_{v}f(\deg_{v})$ gives the theorem. ∎
@@
Using the same idea as in Theorem˜4.3 of collapsing a step from a bad configuration into the next step, and applying Theorem˜2.4 with $k_{1}=n\alpha$ and $k_{2}=\frac{r-1}{rn^{3}\alpha}$, the absorption time is bounded by $O_{r}(n^{4}\alpha^{2})$.
@@ Appendix A Proofs for Preliminaries
Consider the variant of the process where every vertex starts with its own mutant, all with fitness $1$. The process will evolve as usual until a single color takes over the whole graph. A mutant currently occupying subset $S$ fixates if and only if one of the vertices in the subset takes over in the altered process, which in turn equals the fixation probability with only that vertex by viewing all other colors as residents. ∎
@@
We will need the following theorem in [diaz_approximating_2014] to establish Theorem˜2.4.
@@
Now $\phi^{\prime}(Y_{i}^{\prime})$ satisfies the conditions in Theorem˜A.1 with the same parameters, since for all $Y_{i}^{\prime}$ such that $\phi^{\prime}(Y_{i}^{\prime})\notin\{0,k_{1}\}$, the conditions holds by assumption, and if $\phi^{\prime}(Y_{i}^{\prime})=k_{1}$, then $\phi^{\prime}(Y_{i+1}^{\prime})=0$ so the conditions also hold. Applying Theorem˜A.1 gives the claim. ∎
See 2.5
@@
    $\Pr(|\hat{f}-f|>\varepsilon f)\leqslant 2\exp(-2\varepsilon^{2}f^{2}N)\leqslant 2\exp\left(-2\varepsilon^{2}c_{\mathrm{fp}}^{2}n^{-2c_{1}}N\right)\leqslant\tfrac{1}{8},$
@@
    $\Pr(\tau>T)\leqslant\frac{\mathbb{E}[\tau]}{T}\leqslant\frac{C_{\tau}n^{c_{2}}}{8C_{\tau}Nn^{c_{2}}}=\frac{1}{8N}.$

### v2

@@ 1 Introduction
Evolution occurs in populations of reproducing individuals. The modern synthesis in the early $20^{\text{th}}$ century fused natural selection with Mendelian inheritance into a quantitative framework [huxley1942evolution]. Within this framework, population genetics studies how alleles spread through a population, with particular focus on fixation—the event that a mutant allele takes over a resident background [fisher1923xxi, haldane1927mathematical, Wright_1931, moran1959survival, kimura1962probability, kimura1979fixation, kimura1980average, nei1983estimation, charlesworth1997rapid, whitlock2000fixation, przeworski2003estimating, patwa2008fixation, cvijovic2015fate]. Evolutionary graph theory is a general framework to study evolutionary dynamics in structured populations [nowak2003linear, lieberman2005evolutionary, nowak2006evolutionary, ohtsuki_simple_2006]. A common class of models studied in evolutionary graph theory is variations of what are referred to as Moran processes. In this model, a population of size $n$ is represented by a (possibly weighted, directed) graph on $n$ locations, where edges constrain who can replace whom [^1] The original Moran process, described by Patrick Moran in 1958 [enwiki:1277295443, moran1958random], considered a well-mixed population where each individual can replace any other individual, which is equivalent to the graph being complete. Generalizations to the process on graphs were introduced early in the development of evolutionary graph theory [lieberman2005evolutionary, nowak2006evolutionary]. We follow the modern literature (e.g. [diaz_approximating_2014, diaz2016absorption, goldberg2020phase]) and use the term “Moran process” to refer to more generalized processes, where additional features such as graph structure and the order of the birth and death steps may also be considered. [lieberman2005evolutionary]. In this work, we consider unweighted, undirected graphs. Two types of individuals live on the graph, with one individual at each vertex: residents, who have fitness $1$, and mutants, who have fitness $r$.
@@
In this paper, we explore a smooth interpolation between the two processes, which we call the $\lambda$-mixed Moran process. We unify these models with a single parameter $\lambda\in[0,1]$; in each step, the process performs a Bd update with probability $\lambda$ and a dB update with probability $1-\lambda$. While seemingly simple, the model exhibits interesting properties. For example, we give an example where the fixation probability for $\lambda\in(0,1)$ is not bounded by its values at $\lambda=0$ and $\lambda=1$, implying also that monotonicity in $\lambda$ does not hold.
@@ 2.1 Models and Notation
Before delving into specific regimes, we provide two general characteristics for the fixation probabilities in the mixed Moran process. These results were previously known for the pure Bd and dB update rules, and we generalize them to the mixed process. The proofs appear in Section˜A.1.
@@
For any graph $G$, when $r=1$, for all $\lambda$,
@@
Theorem 2.4.
For any graph $G$, any $\lambda$, and any $S\subseteq V$, we have
$r\geqslant 1\implies\mathrm{fp}^{\lambda,r}_{G}(S)\geqslant\mathrm{fp}^{\lambda,1}_{G}(S),\quad r\leqslant 1\implies\mathrm{fp}^{\lambda,r}_{G}(S)\leqslant\mathrm{fp}^{\lambda,1}_{G}(S).$
We believe that one could prove a stronger statement that the fixation probability is monotone in $r$ and the current set $S$. While seemingly intuitive, there is actually some subtlety in proving it rigorously (cf. Remark˜A.1 for more discussions on this). We leave proving the property for future work.
@@ 2.2 Technical tools
We use a potential-function (drift) method. Fix a potential function $\phi$ on subsets $S\subseteq V$. Our goal is to bound the expected one-step change of $\phi$. These theorems are relatively standard, so we defer the proofs to Section˜A.2.
@@ 3 The case $\lambda=1/2$
The claim follows from Theorem˜2.6 and 2.3. ∎
@@
Because $\phi(S)=0$ at $S=\emptyset$ and $\phi(S)=n$ at $S=V$, applying Theorem˜2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ gives
@@
Recall that the fixation probability with initial mutants $S_{0}$ is $|S_{0}|/n$ when $r=1$. Combined with Theorem˜2.4, we yield the following corollary.
@@
Applying Theorem˜2.7 together with Theorem˜3.3 and Corollary˜3.4 gives an FPRAS for Mixed Moran fixation on any graph when $\lambda=1/2$ and $r\geqslant 1$.
@@ 4 Almost Regular Graphs and Random Graphs
We first show that for a regular graph $G$, the fixation probability starting with an initial mutant set $S_{0}$ when $r=1$ is $|S_{0}|/n$, regardless of $\lambda$ and the position of the mutants. This could be compared with Theorem˜3.1 where the same property holds for $\lambda=1/2$ and all graphs: both cases demonstrate this property because the process balances the change in different directions, resulting in the noted invariance. Similar to Theorem˜3.3, we also have the absorption time is $O_{r}(n^{4})$. Since the proofs are very similar, we defer them to Appendix˜B.
@@
For any regular graph $G$, any $\lambda$ and initial mutant set $S_{0}\subseteq V$, we have that $\mathrm{fp}^{\lambda,r=1}_{G}(S_{0})=|S_{0}|/n$ for $r=1$ and the absorption time is at most $O_{r}(n^{4})$ for any $r>0$.
@@
We may assume the graph is not regular, since otherwise the result is already implied by Theorem˜4.1. Thus, $\alpha\geqslant(n-1)/(n-2)$. For $n\geqslant 3$, since $r\geqslant\alpha^{2}$ and $\frac{n-1}{n-2}\geqslant\frac{n}{n-1}$,
@@
Applying Theorem˜2.5 with $k_{1}=n$ and $k_{2}=(r-1)/(rn^{3})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$.
For the fixation probability, applying Theorem˜2.6 with $k_{0}=1$ and $k_{1}=n$ gives $\mathrm{fp}(S_{0})\geqslant\frac{1}{n}$}.
@@
Applying Theorem˜2.5 with $k_{1}=n^{2}$, and $k_{2}=(r-1)/(2rn^{2})$ yields $\mathbb{E}[\tau]$ is $O_{r}(n^{4})$. For the fixation probability, applying Theorem˜2.6 with $k_{0}\geqslant 1$ and $k_{1}=n^{2}$ gives $\mathrm{fp}(S_{0})\geqslant n^{-2}$}. ∎
@@
Combining Theorem˜2.7, 4.3, and 4.5 yields the FPRAS for random graphs.
@@ 5 Bidegreed Graphs
and $\phi(S_{i})$ is a martingale. Applying Theorem˜2.6 with $k_{0}=f(\deg_{u})$ and $k_{1}=\sum_{v}f(\deg_{v})$ gives the theorem. ∎
@@
Using the same idea as in Theorem˜4.3 of collapsing a step from a bad configuration into the next step, and applying Theorem˜2.5 with $k_{1}=n\alpha$ and $k_{2}=\frac{r-1}{rn^{3}\alpha}$, the absorption time is bounded by $O_{r}(n^{4}\alpha^{2})$.
@@ A.1 Basic Properties of mixed Moran Process
Consider a modified process, where instead of two types, there exist $n$ types, all of which have a fitness of $1$. Initially, vertex $i$ starts with type $i$. The process evolves as the $\lambda$-mixed process with fitness $1$:
  • A Bd step happens with probability $\lambda$, selecting a vertex to birth uniformly at random and replacing a random neighbor.
  • A dB step happens with probability $1-\lambda$, selecting a vertex to die uniformly at random and adopting a random neighbor.
The process terminates when the whole graph contains a single type. Recall $\mathrm{fp}(S)$ is the fixation probability in the original process for a mutant currently occupying subset $S$, and we denote by $\widetilde{\mathrm{fp}}(u)$ the probability that type $u$ takes over the whole graph in the modified process. For any vertex $u$, by viewing all types $v\neq u$ as residents and $u$ as the mutant, we have $\widetilde{\mathrm{fp}}(u)=\mathrm{fp}(u)$. By a coupling argument, the probability that the probability of a mutant currently occupying $S$ fixating equals the probability that some $u\in S$ takes over, i.e.,
$\mathrm{fp}(S)=\sum_{u\in S}\widetilde{\mathrm{fp}}(u)=\sum_{u\in S}\mathrm{fp}(u).\qed$
See 2.4
Proof.
We prove the case for $r\geqslant 1$, and the case when $r\leqslant 1$ is implied by swapping the role of mutants and residents and setting fitness of the mutant as $1/r$.
We use a coupling argument. Let $(Z_{t})_{t\geqslant 0}$ denote the neutral process with fitness $1$ and $(X_{t})_{t\geqslant 0}$ denote the advantageous process with fitness $r>1$, both starting from the same set $Z_{0}=X_{0}=S$. We construct a coupling such that $Z_{t}\subseteq X_{t}$ for all steps $t$. Consider the transition from time step $t$ to $t+1$ and any vertex $v\in V$.
  • We bound the probability that a mutant is placed at $v$. Let $p^{Z}_{t+1}(v)$ denote the probability that a mutant is placed at $v$ in $Z_{t+1}$, conditioned on $Z_{t}$, and $p^{X}_{t+1}(v)$ denote the same quantity for process $(X_{t})$. We have
    $p^{Z}_{t+1}(v)$ $=\lambda\sum_{u\in N(v)\cap Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\cap Z_{t}|}{\deg_{v}}$
    $p^{X}_{t+1}(v)$ $=\lambda\sum_{u\in N(v)\cap X_{t}}\frac{r}{w(X_{t})}\cdot\frac{1}{\deg_{u}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{r|N(v)\cap X_{t}|}{w_{v}(X_{t})}$
    Since $r\geqslant 1$ and $Z_{t}\subseteq X_{t}$, we have
    $\sum_{u\in N(v)\cap Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}\leqslant$ $\sum_{u\in N(v)\cap X_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{u}}\leqslant\sum_{u\in N(v)\cap X_{t}}\frac{r}{w(X_{t})}\cdot\frac{1}{\deg_{u}},$
    $\frac{1}{n}\cdot\frac{|N(v)\cap Z_{t}|}{\deg_{v}}\leqslant$ $\frac{1}{n}\cdot\frac{|N(v)\cap X_{t}|}{\deg_{v}}\leqslant\frac{1}{n}\cdot\frac{r|N(v)\cap X_{t}|}{w_{v}(X_{t})},$
    and therefore $p^{Z}_{t+1}(v)\leqslant p^{X}_{t+1}(v)$.
  • We bound the probability that a resident is placed at $v$. Let $q^{Z}_{t+1}(v)$ denote the probability that a resident is placed at $v$ in $Z_{t+1}$, conditioned on $Z_{t}$, and $q^{X}_{t+1}(v)$ denote the same quantity for process $(X_{t})$. Then,
    $q^{Z}_{t+1}(v)$ $=\lambda\sum_{v\in N(v)\setminus Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\setminus Z_{t}|}{\deg_{u}},$
    $q^{X}_{t+1}(v)$ $=\lambda\sum_{v\in N(v)\setminus X_{t}}\frac{1}{w(X_{t})}\cdot\frac{1}{\deg_{v}}+(1-\lambda)\cdot\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{w_{u}(X_{t})}$
    Since $r\geqslant 1$, we have $w(X_{t})\geqslant n$ and $w_{u}(X_{t})\geqslant\deg_{u}$. Furthermore, $N(v)\setminus Z_{t}\supseteq N(v)\setminus X_{t}$. Thus,
    $\sum_{v\in N(v)\setminus Z_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}$ $\geqslant\sum_{v\in N(v)\setminus X_{t}}\frac{1}{n}\cdot\frac{1}{\deg_{v}}\geqslant\sum_{v\in N(v)\setminus X_{t}}\frac{1}{w(X_{t})}\cdot\frac{1}{\deg_{v}},$
    $\frac{1}{n}\cdot\frac{|N(v)\setminus Z_{t}|}{\deg_{u}}$ $\geqslant\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{\deg_{u}}\geqslant\frac{1}{n}\cdot\frac{|N(v)\setminus X_{t}|}{w_{u}(X_{t})},$
    and therefore $q^{Z}_{t+1}(v)\geqslant q^{X}_{t+1}(v)$.
Thus, we can couple the transitions such that if a mutant is placed at vertex $v$ in $(Z_{t})$, a mutant is also placed at $v$ in $(X_{t})$ [^6] Note that it could be the case that $Z_{t}$ or $X_{t}$ already contains $v$. That is fine and does not affect our coupling.; if a resident is placed at vertex $v$ in $(X_{t})$, a resident is also placed at $v$ in $(Z_{t})$. Consequently, $Z_{t}\subseteq X_{t}$ holds for all $t$. We conclude that $\mathrm{fp}^{\lambda,1}_{G}(S)\leqslant\mathrm{fp}^{\lambda,r}_{G}(S)$. ∎
Remark A.1.
We briefly explain the challenge of proving the monotonicity of fixation probability with respect to $r$ in the mixed Moran process.
For pure Bd updates, the challenge in establishing monotonicity in $r$ is that the reproduction probability is normalized by the total fitness, and consequently, having more mutants can potentially slow down the reproduction of other mutants, thereby breaking the natural coupling (see, e.g. [diaz2016absorption] for a more thorough discussion). In [diaz2016absorption], this is overcome by pivoting to a continuous-time Markov chain, where each vertex waits a random time before reproduction, and this time follows an exponential distribution parameterized by its fitness. Then, the discrete sequence of events in this continuous-time Markov chain corresponds exactly to the discrete-time Moran process. Furthermore, in this setting, the presence of additional mutants does not slow down the reproduction of others. Thus, a natural coupling demonstrates stochastic dominance in the fixation probability of the continuous embedding.
However, for the mixed process, this solution faces another challenge: In the continuous embedding for Bd, each vertex has a rate of $f(u)$ for Birth, whereas dB has a fixed rate of $1$ for death. This is problematic since this render will break the condition that the ratio between the rates of Bd and dB events remains $\lambda/(1-\lambda)$; otherwise, the discretization of the continuous-time chain is not equivalent to the discrete Moran process anymore. A natural attempt is then to scale the dB rate by the total fitness to fix this ratio, but this brings back the original problem: having more mutants increases the total fitness and thus the rate of death events, which again breaks the natural coupling. Regardless, we believe that monotonicity in $r$ still holds, although establishing it likely requires a more careful analysis.
A.2 Martingales and Drifts
@@
We now prove the technical lemmas. We will need the following theorem in [diaz_approximating_2014].
@@
Now $\phi^{\prime}(Y_{i}^{\prime})$ satisfies the conditions in Theorem˜A.2 with the same parameters, since for all $Y_{i}^{\prime}$ such that $\phi^{\prime}(Y_{i}^{\prime})\notin\{0,k_{1}\}$, the conditions holds by assumption, and if $\phi^{\prime}(Y_{i}^{\prime})=k_{1}$, then $\phi^{\prime}(Y_{i+1}^{\prime})=0$ so the conditions also hold. Applying Theorem˜A.2 gives the claim. ∎
See 2.6
@@
    $\mathbb{P}(|\hat{f}-f|>\varepsilon f)\leqslant 2\exp(-2\varepsilon^{2}f^{2}N)\leqslant 2\exp\left(-2\varepsilon^{2}c_{\mathrm{fp}}^{2}n^{-2c_{1}}N\right)\leqslant\tfrac{1}{8},$
@@
    $\mathbb{P}(\tau>T)\leqslant\frac{\mathbb{E}[\tau]}{T}\leqslant\frac{C_{\tau}n^{c_{2}}}{8C_{\tau}Nn^{c_{2}}}=\frac{1}{8N}.$
@@
Appendix B Regular Graphs
See 4.1
Proof.
We first establish the fixation probability when $r=1$. Fix $S_{0}=\{u\}$ for an arbitrary vertex $u$. Consider the potential $\phi(S)=|S|$. Since all degrees are equal,
Thus, $\phi(S)$ is a martingale. The theorem follows from Theorem˜2.6 and 2.3.
We now bound the absorption time. Let the graph $G$ be $d$-regular. Using the exact same argument as in the proof of Theorem˜3.1, we obtain that the absorption time is $O(n^{3}d)$ when $r=1$. It remains to show the case where $r\neq 1$. We establish the result for $r>1$. The case $r<1$ is symmetric by swapping the roles of mutants and residents.
Consider a boundary edge $(u,v)\in\mathrm{bdry}(S)$. The expected drift contributed by this edge is:
$\lambda\cdot\psi_{u,v}^{Bd}(S)+(1-\lambda)\cdot\psi_{u,v}^{dB}(S)$ $=\frac{\lambda}{w(S)}\left(\frac{r}{d}-\frac{1}{d}\right)+\frac{1-\lambda}{n}\left(\frac{r}{w_{v}(S)}-\frac{1}{w_{u}(S)}\right)$
The Bd component is strictly positive. Since $w(S)\leqslant nr$, we have:
$\frac{\lambda}{w(S)}\left(\frac{r}{d}-\frac{1}{d}\right)\geqslant\frac{\lambda(r-1)}{rnd}.$
For the dB component, we perform an analysis similar to that in the proof of Theorem˜4.7.
  • If all neighbors of the mutant $u$ are residents and all neighbors of the resident $v$ are mutants, then $w_{u}(S)=d$ and $w_{v}(S)=rd$. The dB term vanishes:
    $\frac{r}{w_{v}(S)}-\frac{1}{w_{u}(S)}=\frac{r}{rd}-\frac{1}{d}=0.$
  • If $u$ has at least one mutant neighbor, then $w_{u}(S)\geqslant r+(d-1)$. Thus:
    $\frac{r}{w_{v}(S)}-\frac{1}{w_{u}(S)}\geqslant\frac{r}{rd}-\frac{1}{d+r-1}=\frac{1}{d}-\frac{1}{d+r-1}\geqslant\frac{r-1}{d(d+r)}.$
  • If $v$ has at least one resident neighbor, then $w_{v}(S)\leqslant r(d-1)+1$. Thus:
    $\frac{r}{w_{v}(S)}-\frac{1}{w_{u}(S)}\geqslant\frac{r}{r(d-1)+1}-\frac{1}{d}\geqslant\frac{r-1}{d^{2}r}.$
We call a configuration “bad” if every mutant has only resident neighbors and every resident has only mutant neighbors.
In a bad state, the drift is non-negative, and a bad state converts into a non-bad state in one step. Furthermore, in a non-bad state, we have
$\mathbb{E}[\phi(S_{i+1})-\phi(S_{i})\mid S_{i}]\geqslant\frac{r-1}{rn^{2}d}.$
Consider a modified Markov chain $S^{\prime}$ which “glues” one step in a bad configuration with one more step, which preserves the absorption time up to a constant factor and satisfies
$\mathbb{E}[\phi(S_{i+1}^{\prime})-\phi(S_{i}^{\prime})\mid S_{i}]\geqslant\frac{r-1}{rn^{2}d}.$
Since $\phi(V)=n$, applying Theorem˜2.5 yields an absorption time of $\frac{r}{r-1}O(n^{3}d)=O_{r}(n^{4})$. ∎